The Population of Japan is 127 million (2018).

The Population of South Korea is 50.6 million in Sep 2016.

90% are Lactose Intolerant.

1. How many Heterozygous people are in Japan? (Heterozygous for Lactose Intolerance).
Answer #1

Answer let homozygous recessive responsible for lactose intolerance in person..

So the population frequency is 90%. This means that q2 (percentage of individual inherited trait)

So q2 = 90% or 0.9

then q ( allelic frequency in population) = root of 0.9 = 0.95

then p = 1-0.95 = 0.05

so heterozygous =2pq then 2(0.95×0.05) = 0.095

hence 0.095 people are heterozygous in Japan population